OpenStudy (anonymous):

-5x+8y=42 find the value of y when x equals 6

OpenStudy (anonymous):

can anybody solve this problem

OpenStudy (amoodarya):

-5x+8y=42 when x=6 put in equation -5(6)+8y=42---->8y=42+30 8y=72 y=9